### v3

# Let’s Fuse Step by Step: A Generative Fusion Decoding Algorithm with LLMs for Multi-modal Text Recognition

Abstract
We introduce “Generative Fusion Decoding” (GFD), a novel shallow fusion framework, utilized to integrate Large Language Models (LLMs) into multi-modal text recognition systems such as automatic speech recognition (ASR) and optical character recognition (OCR). We derive the formulas necessary to enable GFD to operate across mismatched token spaces of different models by mapping text token space to byte token space, enabling seamless fusion during the decoding process. The framework is plug-and-play, compatible with various auto-regressive models, and does not require re-training for feature alignment, thus overcoming limitations of previous fusion techniques. We highlight three main advantages of GFD: First, by simplifying the complexity of aligning different model sample spaces, GFD allows LLMs to correct errors in tandem with the recognition model, reducing computation latencies. Second, the in-context learning ability of LLMs is fully capitalized by GFD, increasing robustness in long-form speech recognition and instruction aware speech recognition. Third, GFD enables fusing recognition models deficient in Chinese text recognition with LLMs extensively trained on Chinese. Our evaluation demonstrates that GFD significantly improves performance in ASR and OCR tasks, with ASR reaching state-of-the-art in the NTUML2021 benchmark. GFD provides a significant step forward in model integration, offering a unified solution that could be widely applicable to leveraging existing pre-trained models through step by step fusion.  [^1] Code is available at https://github.com/mtkresearch/generative-fusion-decoding
1 Introduction
In recent years, integrating Large Language Models (LLMs) into multi-modal systems has emerged as a frontier, offering substantial advancements in applications such as automatic speech recognition (ASR) [40], visual question answering (VQA) [36], reinforcement learning [50], and generalist agent [42]. The robust language model capabilities that the LLM offers make it an ideal candidate for integration with other models, many of which are “undertrained” in the textual domain, compared to trillions of tokens consumed by the LLMs. Obtaining sufficient quality paired data and training resources in a huge endeavor for text recognition systems such as automatic speech recognition (ASR) and optical character recognition (OCR).
The integration of heterogeneous models has been explored extensively in ASR literature, as exemplified by strategies such as shallow fusion, late fusion, mid fusion, and early fusion. Works in shallow fusion [9, 27, 10] interpolates text recognition outputs with LM outputs at each step of beam search. Late fusion approaches typically engage LLMs to score the initial sentence generated by the recognition model. The scores are then utilized for fusing with recognition probabilities for joint transcription [6], for re-ranking the the constrained set of hypothesis as work in re-scoring [46], for generating new hypothesis options through Generative Error Correction (GER) [5]. Notably, methods in mid-fusion propose fusing models through model attentions to integrate LLM into the decoding process of the recognition model [41, 37], and work in early-fusion has proposed concatenating speech embeddings or image embeddings along with LLM embedings as inputs to the LLM for ASR and VQA [17, 7]. However, the aforementioned fusion methods bear several challenges. The early fusion techniques are often limited by a low correction rate due to the restricted range of available choices for re-evaluation. Mid-fusion discards the whisper decoder, which has been shown to have denoising ability[19]. Additionally, late fusion with GER may lead to hallucination due to the LLM as the LLM generation is devoid of auditory inputs. Due to the requirements for shallow and late fusion to align output spaces, and the need for mid and early fusions to involve joint training with additional paired data, prevent an LLM from being used as is with existing text-recognition models without proper feature alignment through trainings.
Figure 1: The GFD integrated framework. The framework aims to integrate pre-trained text-recognition models with LLMs to augment the recognition capabilities. However, direct fusion is impeded by the discrepancy in token spaces between the two model types. To overcome this obstacle, GFD is introduced, which converts the token space into a byte space, thereby facilitating incremental fusion. Equations 7 and 8 are employed to compute the fusion probability, enabling a stepwise fusion process. The symbols $\mathcal{Z}^{\texttt{TR}}$ and $\mathcal{Z}^{\texttt{LM}}$ denote the contextual information pertinent to the text-recognition models and the language models, respectively. Notably, $\mathcal{Z}^{\texttt{TR}}$ may correspond to either an audio input for ASR or a text image for OCR, whereas $\mathcal{Z}^{\texttt{LM}}$ typically represents a textual prompt.
To address the aforementioned challenges, we introduce a novel shallow fusion framework, termed “Generative Fusion Decoding” (GFD). GFD resolves the token space mismatch between models by bridging heterogeneous spaces through the expected score values of plausible sequences and mapping text token space to byte token space. This allows GFD to seamlessly fuse LLMs with text recognition models during the decoding process. Our unified framework is plug-and-play, applicable to text recognition tasks such as ASR and OCR, and compatible with diverse auto-regressive text recognition models as well as language models. We highlight three main advantages of GFD. First, GFD addresses key limitation of previous fusion techniques — namely the complexity of aligning different model sample spaces via re-training associated with prior deep integration [41] or two-pass [46] strategies. It allows LLMs to participate in the sampling process as conventional shallow fusion, providing real-time correction of errors generated by the text recognition model, thereby extending the influence of LLMs beyond the complete outputs of the recognition model. Secondly, as modern LLMs are trained on trillions of multilingual text tokens [22, 26], generally around 3 orders of magnitude more than end-to-end ASR or OCR models [40], LLMs in GFD can be seen as a strong scoring model in the step by step decoding that improves the multi-lingual outputs of ASR and OCR systems. Lastly, because of the in-context learning ability of LLMs, the GFD fusion system with LLMs and ASR can readily leverage domain information as prompts effectively for domain adaptation similarly to prior work through re-scoring [14] or instruction tuning [35].
To evaluate the efficacy of our algorithm, we implemented the GFD for ASR and OCR tasks. In the ASR task, we evaluated on the conventional ASR task as shown in Section 4.2 for English and Chinese audios. The evaluation of Chinese ASR is motivated by existing literature [40], which indicates a performance outlier in Chinese compared to other languages despite abundant paired-training data. We hypothesized that LLMs could enhance Chinese ASR by offering additional scoring to differentiate between semantically distinct but homophonous words. Furthermore, we demonstrate that ASR with GFD can improve recognition for long-form audio through leveraging the long-context ability of LLMs (Section 4.3); for context-aware recognition for rare words and domain tags via in-context learning capabilities of LLMs (Section 4.4). Lastly, application to the OCR task, particularly targeting long text OCR for historical documents, demonstrates that OCR with GFD can significantly improve recognition accuracy (Section 4.5).
2 Related work
2.1 Model fusion
Training a multi-objective model from scratch is often costly [3, 2, 1, 15]. Consequently, researchers have pivoted towards combining existing models with different modalities to improve accuracy without the prohibitive costs of building new systems from the ground up.
Model fusion developed in the field of ASR provides a plausible path for combining existing trained models. The technique have evolved significantly in recent years, encompassing a variety of approaches designed to integrate different models to enhance performance. Shallow fusion or late fusion, often employed in ASR, combines end-to-end ASR models with external language models at the decoding level, improving recognition accuracy without altering the underlying ASR architecture [27, 25, 6, 52]. However, due to the heterogeneous sample spaces of models, the prerequisite of shallow and late fusion requires aligning sample spaces of model distributions, enabled through fine-tuning a projection module [6]. Late fusion training methods may suffer from modality laziness problem in tasks where uni-modal priors are meaningful [16]. Work in re-scoring N-best hypothesis from the first pass inference with an external language model has been shown effective in reducing recognition errors [43, 23, 46]. However, the method prunes the correct prefix early on away from N-best list and requires a second pass in the decoding process.
Alternatively, deep fusion integrates models at the level of hidden feature, requiring fine-tuning models to fuse deep features [20]. Cross-modal fusion, similar to deep fusion, integrates pre-trained end-to-end ASR model with LLM [41, 51, 33] or vision model with LLM [7, 36] via learning a joint representation with large amount of extra paired audio-text or image-text data.
Our newly proposed approach—GFD—decodes in homogeneous sequence element space with byte strings, and therefore relaxes the alignment of sample spaces of distributions without any model training. Without pre-training or fine-tuning, we ensure that the individual pre-trained model performance is guaranteed on their respective task, and is not influenced by volatility of further training.
2.2 Contextual conditioning
Auto-regressive LLMs have exhibited capabilities in in-context learning [39], instruction following [38], and knowledge synthesis [44]. Such capabilities have been applied to solving domain adaptation in speech recognition for rare words or out-of-domain context through contextual biasing [10] and prompting fine-tuned models [35, 47, 33, 47]. With GFD, such problem setting can be tackled by leveraging a performant LLM as is through prompting without further fine-tuning.
2.3 Chinese ASR
One of the most significant challenges in developing ASR systems for the Chinese language stems from its highly homophonous nature [30, 29, 8]. Unlike English, where there is a larger variety of phonemes and a relatively consistent correspondence between spelling and sound, Chinese relies on a limited set of tones and syllables to represent thousands of characters. Consequently, Chinese ASR systems must not only accurately capture the tonal nuances but also analyze the linguistic context to disambiguate these homophones. This requires sophisticated LMs that can effectively leverage contextual clues to predict the intended characters. The integration of LLMs has shown promise in addressing these challenges [11, 31, 34]. GFD is also expected to enhance the performance of Chinese ASR by utilizing the contextual conditioning capabilities of LLMs.
3 Method
3.1 Generative fusion decoding
For conditional text generation models, the sequence with the highest probability during inference is found using the following formula:
$\{T_{s}\}^{*}=\underset{\{T_{s}\}}{\mathrm{arg\,max}}\log\mathbb{P}(\{T_{s}\},\mathcal{Z}),$ (1)
where $\{T_{s}\}$ represents the sequence of tokens generated by the model, and $\mathcal{Z}$ represents the given context or conditioning information, such as audio for speech recognition models [40], images for vision-language-models [1], and prompts for typical language models [4].
Auto-regressive generation is one approach to realize conditional text generation. In this approach, the auto-regressive model is conditioned on the previously generated tokens to generate the next token sequentially. Therefore, the probability $\log\mathbb{P}(\{T_{s}\}|\mathcal{Z})$ is typically decomposed using the chain rule of probability, as follows:
$\log\mathbb{P}(\{T_{s}\},\mathcal{Z})=\sum_{s=1}^{S}\log\mathbb{P}(T_{s}\mid T_{<s},\mathcal{Z}),$ (2)
where $T_{s}$ is the token at position $s$ in the sequence, and $T_{<s}$ represents all the tokens preceding position $s$. In real-world applications, it is impracticable to enumerate all possible token sequences, so beam search is typically used as an approximation method.
In the setting of shallow fusion, we attempt to let multiple models jointly determine the sequence, as shown in the following formula rewrote from Equation (1):
$\{T_{s}^{\texttt{fuse}}\}^{*}$ $=\underset{\{T_{s}^{\texttt{fuse}}\}}{\mathrm{arg\,max}}\sum_{m}\lambda_{m}\log\mathbb{P}_{m}(\{T_{s}^{(m)}\}=\{T_{s}^{\texttt{fuse}}\},\mathcal{Z}^{(m)})$ (3)
where $\{T_{s}^{\texttt{fuse}}\}$ represents the fused sequence of tokens generated by combining the outputs of multiple models, $\lambda_{m}$ is a weighting factor for the $m$-th model, $p_{m}$ denotes the probability distribution of the $m$-th model and $\mathcal{Z}^{(m)}$ represents the context or conditioning information specific to the $m$-th model.
When the sample spaces of models are the same, Equations (3) and (2) can be combined to realize incremental fusion. If the models have different sample spaces due to a mismatch in token spaces, there is no simple way to achieve incremental fusion. One alternative method to approximate Equation (3) is to fuse at the level of the fully generated results from each model. Nevertheless, in practice, fusion at the level of fully generated results poses the problem of an enormous search space because different conditioning variables $\mathcal{Z}^{(m)}$ may produce vastly different results.
To address these challenges, we have introduced a probability transformation, denoted as $\mathcal{M}^{(m)}$, that converts token-level representations into byte-level representations:
$\mathcal{M}^{(m)}:\mathbb{P}_{m}(\{T_{s}\},\mathcal{Z}^{(m)})\longrightarrow\mathbb{P}_{m}(\{B_{l}\},\mathcal{Z}^{(m)}),$ (4)
where $\{B_{l}\}$ represents the sequence of bytes after the transformation, and $l$ denotes the position in the byte sequence. This transformation allows for a unified representation across different models, facilitating the fusion process even when the original token spaces differ. The byte-level fusion can then be performed using a similar approach to Equation (3), but with the byte-level probabilities:
$\{B_{l}^{\texttt{fuse}}\}^{*}$ $=\underset{\{B_{l}^{\texttt{fuse}}\}}{\mathrm{arg\,max}}\sum_{m}\lambda_{m}\log\mathbb{P}_{m}(\{B_{l}\}=\{B_{l}^{\texttt{fuse}}\},\mathcal{Z}^{(m)}).$ (5)
To operationalize the probability transformation $\mathcal{M}^{(m)}$, we define a mapping from the token-level probabilities to the byte-level probabilities. This mapping takes into account the prefix relationship between the token sequence and the byte sequence. Specifically, we express the byte-level probability $\mathbb{P}_{m}(\{B_{l}\},\mathcal{Z}^{(m)})$ as a sum over all possible token sequences that share a common prefix with the byte sequence $\{B_{l}\}$. The probability of each token sequence is computed as the product of the conditional probabilities of each token given the preceding tokens and the context $\mathcal{Z}^{(m)}$. This relationship is formalized in the following equation:
$\mathbb{P}_{m}(\{B_{l}\},\mathcal{Z}^{(m)})=\sum_{\{B_{l}\}}\biggr{[}\prod_{s}\mathbb{P}_{m}(T_{s}^{(m)}\mid T_{<s}^{(m)},\mathcal{Z}^{(m)})\biggr{]}_{\{T_{s}\}}\times\mathbb{1}\bigl{(}\{T_{s}\}\texttt{.prefix}=\{B_{l}\}\bigl{)},$ (6)
where .prefix is a function that maps a token sequence to its corresponding byte sequence prefix, and $\mathbb{1}$ is the indicator function that is 1 if the token sequence has the specified byte sequence as a prefix and 0 otherwise. The indicator function ensures that only those token sequences that can be mapped to the given byte sequence contribute to the byte-level probability. In Equation (6), the complexity remains high at $\mathcal{O}(V^{S})$, where $V$ represents the vocabulary size of tokens and $S$ is the sequence length, for identifying token sequences that match the specified criteria of the indicator.
Figure 2: Example of the main sequence and alternative tokens. Assume that the byte sequence is "abcdefghijk". The main token sequence is the tokenization result of the byte sequence and is denoted as $\{\hat{T}_{s}\}$. The alternative tokens are denoted as $T^{\prime}_{i}$.
To reduce complexity, we posit that the main token sequence and its branching alternatives, as shown in Figure 2, dominate the probability contribution, where the main token sequence is produced by applying model tokenization on the byte string and the alternative tokens are essentially look-aheads for potential main tokens that may emerge as the decoding progresses. The significance of the main token sequence is justified by its alignment with the model’s inputs during the original pretraining phase; any other slicing method is penalized in terms of probability due to its lack of representation in the training data. Based on this assumption, we can narrow down the search for token sequences that meet the criteria of $\{T_{s}\}\texttt{.prefix}=\{B_{l}\}$ to only the main token sequence and its branching alternative tokens. We define the main token sequence as $\{\hat{T}_{s}\}$. Given this simplification, we can approximate the byte-level probability $\mathbb{P}_{m}(\{B_{l}\},\mathcal{Z}^{(m)})$ by considering only the main token sequence and its immediate alternatives that share the same prefix with the byte sequence $\{B_{l}\}$. This approximation significantly reduces the computational complexity to $\mathcal{O}(V\times S)$ and is expressed in the following equation:
$\mathbb{P}_{m}(\{B_{l}\},\mathcal{Z}^{(m)})\approx$  $\texttt{~{}}\mathbb{P}_{m}(\hat{T}_{1}\mid\hat{T}_{<1},\mathcal{Z}^{(m)})\times\bigl{[}$ (7)
 $\texttt{~{}~{}~{}}\mathbb{P}_{m}(\hat{T}_{2}\mid\hat{T}_{<2},\mathcal{Z}^{(m)})\times\bigl{[}$
[…]
 $\texttt{~{}~{}~{}~{}~{}~{}~{}}\mathbb{P}_{m}(\hat{T}_{S}\mid\hat{T}_{<S},\mathcal{Z}^{(m)})$
 $\texttt{~{}~{}~{}~{}~{}~{}~{}}+\sum_{t}\mathbb{P}_{m}(t\mid\hat{T}_{<S},\mathcal{Z}^{(m)})\cdot\mathbb{1}\bigl{(}\{\hat{T}_{<S},t\}\texttt{.prefix}=\{B_{l}\}\bigl{)}$
[…]
 $\texttt{~{}~{}~{}}\bigl{]}+\sum_{t}\mathbb{P}_{m}(t\mid\hat{T}_{<2},\mathcal{Z}^{(m)})\cdot\mathbb{1}\bigl{(}\{\hat{T}_{<2},t\}\texttt{.prefix}=\{B_{l}\}\bigl{)}$
 $\texttt{~{}}\bigl{]}+\sum_{t}\mathbb{P}_{m}(t\mid\hat{T}_{<1},\mathcal{Z}^{(m)})\cdot\mathbb{1}\bigl{(}\{\hat{T}_{<1},t\}\texttt{.prefix}=\{B_{l}\}\bigl{)},$
where $t$ represents an alternative token that could follow the sequence $\hat{T}_{<s}$ and meet the criteria of $\{T_{s}\}\texttt{.prefix}=\{B_{l}\}$. Eventually, substituting Equation (7) into (5) can realize generative fusion decoding (GFD).
In summary, our proposed method for conditional text generation, GFD, through late fusion and byte-level probability transformation offers a novel way to integrate the outputs of multiple models with different token spaces. By transforming token-level probabilities to byte-level probabilities and focusing on the most probable token sequences, we can efficiently fuse model outputs.
3.2 Fusing text-recognition models with LLM
To evaluate the efficacy of our algorithm, we implemented GFD for ASR and OCR tasks. This is achieved by fusing pre-trained text-recognition models with LLMs to enhance recognition capability. In such scenarios, the proposed tokens of ASR and OCR model are conditioned on a modal signal $\mathcal{Z}^{\texttt{TR}}$ while that of the LLM is not. Therefore, the same next-byte probability estimation from the LLM and text recognition model might stem from different token choices. Although the difference would converge eventually as the sequence rolls out, given limited beams in the beam search process, the potential overestimation of implausible words due to scores contributed by LLM spikes the probabilities of common word-leading bytes such as ’t’ and ’r’, and produces locally unstable rankings that may result in false eliminations of correct beams in decoding steps. In response to this challenge, we have adopted text recognition models (ASR and OCR) to propose the next byte/token, while employing LLMs to provide delayed corrective feedback on past tokens. The fusion decoding methodology we have adopted, derived from Equation (5), is delineated in the following formula:
$\{B_{1},..,B_{t}\}^{*}=\underset{\{B_{1},..,B_{t}\}}{\mathrm{arg\,max}}\bigl{[}(1-r)\cdot\log\mathbb{P}^{\texttt{TR}}(\{B_{1},..,B_{t}\},\mathcal{Z}^{\texttt{TR}})$ (8)
$+r\cdot\log\mathbb{P}^{\texttt{LM}}(\{B_{1},..,B_{t-k}\},\mathcal{Z}^{\texttt{LM}})\bigl{]},$
where $\{B_{1},\ldots,B_{t}\}^{*}$ represents the optimal sequence of bytes up to and including position $t$, $p^{\texttt{TR}}$ and $p^{\texttt{LM}}$ are the probability distributions of the text-recognition models and the language models, respectively, $\mathcal{Z}^{\texttt{TR}}$ and $\mathcal{Z}^{\texttt{LM}}$ are the contextual information for the text-recognition and language models, respectively, $r$ is a weighting factor that balances the influence of the text-recognition models and the language models, and $k$ depends on the length of the byte/token proposal. As shown in Figure 1, $\mathcal{Z}^{\texttt{TR}}$ can be either an audio or a text image, and $\mathcal{Z}^{\texttt{LM}}$ can be a prompt.
4 Experiment
4.1 Experimental setup
We evaluate the application of GFD to ASR and OCR tasks. For ASR task, we benchmark datasets in English, Chinese, and Cantonese. The deliberate choice of Chinese and Cantonese is due to that the languages are homophonous and tonal, challenging for ASR systems. This complexity is corroborated by the Fleurs experiment in Whisper [40], where the Chinese word error rate is well above the regressed word error rate in comparison to evaluations in other language at the same amount of pre-training data. Of all tested languages in Whisper, it is the only large-scale language (> 10k hours audio) with such a phenomenon. For OCR, we benchmark image dataset containing long sequence of text as we hypothesize that LLM provide semantic information to an OCR model of recognizing long text sequences.
In terms of the evaluated models, we choose Whisper-large-v2 as the ASR model for greedy and beam search. For our proposed GFD, we use Mistral [26] for English datasets and Breeze [22] for Chinese and Cantonese datasets. We term the GFD fused model with Whisper and Mistral as GFD-ASR-EN; that with Whisper and Mistral as GFD-ASR-ZH. In addition, we benchmark GER, based on task-activating prompting method [5, 48], with Mistral-Instruct and Breeze-Instruct models without finetuning. We include two oracle word error rates following previous work[24], where the N-Best Oracle $o_{nb}$ denotes the error rate calculated with the best candidate in the N-best list, and the Compositional Oracle $o_{cp}$ is the best achievable word error rate using all tokens N-best the list. As to OCR, we utilize TrOCR [32], a pre-trained OCR model, and denote the fused model GFD-OCR.
We benchmark models on various datasets, including Librispeech [21], Medical [18], ATCO2 [45], Fleurs [12], NTUML2021[49], and FormosaSpeech  [^2] https://huggingface.co/datasets/MediaTek-Research/formosaspeech for ASR; NAF [13] for OCR. Librispeech is a collection of corpus from audiobooks with subsets Clean and Other. The Librispeech-Noise is its noised variant, featuring various signal-to-noise ratios, ideally for testing ASR system’s robustness to noise. Medical dataset contains 8.5 hours of medical conversations with associated symptom tags to each audio-text pairs. ATCO2 contains audios of air traffic control communication and accompanied meta-information of airports. Fleurs is a multilingual speech corpus. We deliberately choose Cantonese subset for evaluation as the language is homophonous and tonal. NTUML2021 corpus consists of lecture recordings from the “Machine Learning” course at National Taiwan University in 2021, with corresponding transcriptions and English translations labeled by over 20 bilingual native Chinese speakers. FormosaSpeech corpus includes Chinese recordings of Taiwanese accents amassing up to 6.4 hours of audio-text pairs. NAF consists of images from U.S national archives with labelled bounding boxes and annotations, ideal for evaluating OCR performance.
For all ASR experiments, we report word error rate (WER) for evaluations on English datasets, and mix error rate (MER) for that on Chinese datasets. For OCR experiments, we report character error rate (CER) and exact match (EM). For exact match, we require that the recognized texts fully matches the ground truth. All GFD fused models can be run on a single A6000 GPU. We fix $r=0.2$ in Equation 8 for all our experiments; number of beams equaling 5 and 10 for ASR and OCR experiments, respectively.
4.2 Conventional speech recognition
We first verify the efficacy of GFD in conventional speech recognition setting and report results in Table 1. We notice that beam search improves consistently upon greedy search, and use beam search as the baseline. We found that GFD moderately improves on the Medical and Fleurs-HK dataset. In the more challenging NTUML2021, we obtained a state-of-the-art 8.83 mixed error rate, surpassing prior work[49] and even the oracle N-Best score. This improvement is mostly preserved even when we narrow the target space to only consider English terms in the dataset (Table 2), making GFD-ASR an elegant solution for code-switched ASR. Examples in the Appendix show that improvement can be attributed to the LLM correcting grammatical mistakes and domain terminologies. Performance across Librispeech reveals that GFD improves most on moderate noise conditions but diminishes when the noise is too large $(S/R=-5)$.
In addition, we do not find improvements in the GER setting using non-GER-specific instruct models, consistent with previous work[5]. As we already patched malformatted outputs with the best beam result, this increase of error rate is mostly attributed to LLM hallucinations, including incorrect deletions. We therefore posit that GFD is more robust for incorporating off-the-shelf LLMs on the ASR task than GER.
Table 1: Results on conventional speech recognition.
Whisper GER GFD-ASR-EN Oracle
EN dataset Greedy 5-beams (Baseline) $\text{O}_{\text{nb}}$ $\text{O}_{\text{cp}}$
Librispeech-Clean 2.30 2.28 3.21 2.29 1.91 1.63
Librispeech-Other 5.23 4.97 6.57 4.99 4.20 3.43
Medical 7.30 7.22 7.49 6.74 6.17 5.05
Librispeech-Noise ($S/R=10$) 3.50 3.12 4.05 3.02 2.32 1.94
Librispeech-Noise ($S/R=5$) 5.67 5.25 6.00 4.96 4.10 3.28
Librispeech-Noise ($S/R=0$) 15.23 13.54 14.44 13.27 11.66 9.04
Librispeech-Noise ($S/R=-5$) 49.09 47.05 47.35 46.98 43.62 33.19
Whisper GER GFD-ASR-ZH Oracle
ZH or HK dataset Greedy 5-beams (Baseline) $\text{O}_{\text{nb}}$ $\text{O}_{\text{cp}}$
Fleurs-HK (Cantonese) 7.49 6.88 7.33 6.23 5.58 4.58
NTUML2021 11.11 9.97 9.87 8.83 8.88 4.54
NTUML2021 FormosaSpeech
Full English
Whisper (short) 9.97 15.21 -
SICL (short) [49] 9.5 - -
GFD-ASR-ZH (short) 8.83 13.51 -
Whisper (long) 8.40 11.56 27.41
GFD-ASR-ZH (long) 8.18 11.43 26.72
Table 2: Comparison results on long-form and short-form.
4.3 Long-form speech recognition
As LLMs are capable in attending to long context of sequence and possess multilingual capabilities, we posit that GFD-ASR-ZH can improve on long-form audio transcription. We therefore evaluate the long-form transcription performance on NTUML2021 and FormosaSpeech. In constrast to the short-form evaluations in Table 1 in NTUML2021, we prepend all previous transcription from contiguous clips in the same lecture as prompts for both ASR and LLM to realize long form transcriptions, but truncate it when necessary for whisper due to context length limitations. In the setting of evaluations on contiguous short clips, GFD-ASR-ZH (long) consistently outperforms Whisper (long) in both NTUML2021 and FormosaSpeech, demonstrating that the long-context capability of LLM can be effectively utilized through GFD.
4.4 Context-aware speech recognition
Table 3: Results on context-aware ASR task. We set num beams=5 across all our experiments.
Method Prompt ATCO2 Librispeech FormosaSpeech Medical
ASR LLM Norm Raw Clean Other
Whisper No - 47.70 66.44 2.28 4.97 27.42 7.22
Whisper Yes - 31.34 42.37 - - - 6.24
Clairaudience[35] Yes - 28.77 - - - - 6.54
GFD-ASR-EN No Yes 38.75 52.24 2.20 4.61 26.29 6.62
GFD-ASR-EN Yes Yes 25.79 32.46 - - - 6.26
In context-aware speech recognition, we aim to investigate the capability of the fusion model in leveraging contextual information, which is highly desired in real-life contexts, where the incoming speech is expected to be bounded by domain, contain certain specific words, or follow a certain transcribing format. We utilize prompting with ASR model and the LLM model (in the GFD setting) to condition on these external hints, in three following settings.
Domain and Rare Word Prompting: We tested domain-conditioned ASR on the Medical dataset, where the symptom tags are provided along with the speech content. Results show that our GFD method with LLM prompting improves on GFD without prompting, showing prompt sensitiveness of the LLM in the GFD system. However, we did not find further improvement upon whisper prompting, compared to double-prompted GFD. For verifying rare word prompting capabilities, we used the augmented Librispeech and FormosaSpeech dataset, where a target rare word is mixed with 100 other distractors for each data point [28]. For the FormosaSpeech Dataset, we created the rare words with ChatGPT, and generate distractors with a similar approach using the training set. At this scale of distractors, it is unrealistic to prompt on Whisper, as the remaining context length is often insufficient for ASR decoding. By prompting on the LLM using GFD, we demonstrated up to 6% WERR over non-prompting methods on the Librispeech dataset, and 1.6% WERR on the FormosaSpeech Dataset.
Instruction Aware Prompting: We evaluated formatted speech recognition on the ATCO2 dataset, a dataset on air traffic control communications, which has strict regulations on call signs and transcribe formats. By incorporating a LLM, we are able to prompt it with over 4000 words of guidelines from an enitre instruction manual [^3] https://www.faa.gov/air_traffic/publications/atpubs/aim_html/chap4_section_2.html. a possibility not present with Whisper. For Whisper prompting, we include all special call signs and three example sentences extracted from the manual, where we leave details in the Appendix. Results in Table 3shows that GFD with ASR and LLM prompting obtains best results with a WERR of 17.7% compared with Whisper in the normalized (ATCO2-Norm) setting, even outperforming Clairaudience[35], a fine-tuned prompt conditioning model. We also reported scores without word conversion normalization (ATCO2-Raw), from observations that standard normalization can excessively correct errors that conflicts with the transcribing guidelines. In this setting, our gains are even larger, showing instruction following abilites of the GFD system.
4.5 Optical Character Recognition
Table 4: Evaluation on OCR task
CER $\downarrow$ Exact Match $\uparrow$
TrOCR 12.02 24.14
GFD-OCR 10.55 33.33
We demonstrate that GFD can be applied to fusing end-to-end OCR model with LLMs, using TrOCR and Mistral [32, 26]. We experiment the fusion model on a dataset extracted from National Archive Forms dataset [13], denoted NAF-Long. The NAF-Long dataset consists of images with text sequences having character lengths between 100 and 200, equivalent to around 25 to 50 text tokens. In Table 4, we show that fusing LLM to the OCR model significantly improves the OCR results by 16.7 % in character error rate reduction and 38.07 % in exact match improvement.
5 Discussion
Further comparisons with GER. The GFD algorithm relates to GER in that both algorithms perform the selection of output sequences with beam decoding. However, they differ in compute execution and the diversity of sample sequences. First, in GFD, the LLM works in parallel with the ASR decoder, and thus the computation of per step inference can be executed asynchronously with a load bounded by $\mathcal{O}(Z)+\mathcal{O}(k\cdot\max(S_{ASR},{S_{LLM}}))$, where $Z$ denotes the size of speech encoding, $k$ is the beam size, and $S_{ASR}$ and $S_{LLM}$ are ASR and LLM decoding lengths, respectively. The LLM decoding complexity expression matches that of Section 3.2, treating the vocabulary size as a constant in this dicussion. In contrast, GER operates sequentially, requiring the completion of beam decoding with ASR prior to a correction with LLM. This means the execution time of GER is bounded by $\mathcal{O}(Z)+\mathcal{O}((k+1)\cdot S_{LLM})+\mathcal{O}(k\cdot S_{ASR})$, where the additional $O(S_{LLM})$ comes from LLM decoding. Secondly, in GFD, the searched token space attended by the LLM is at least n_beams * sequence length, whereas in GER, the top-k cutoff of the ASR step does not promote diversity between the candidates, which limits the LLM search space. Aside from these differences, GFD and GER are methodologically orthogonal, allowing for their combination in the pursuit of further improvement.
Limitations. Despite promsing results, we identified three cases where GFD is most susceptible to failure, which is an exciting future direction. We categorize them into ASR errors and LLM errors.
ASR proposed errors: When ASR proposes a candidate with high probability that deviates from the phonetic constraint, the LLM may falsely pickup the sequence, inducing an error. There are two main types - semantically activated tokens and time delayed activated tokens. Semantically activated tokens are tokens that are encoded in similar output embedding space due to semantic similarity. We found that these errors are much more common in Chinese and identified some pairs from a top-k analysis of a decoding sequence, e.g., {CJK}UTF8bsmi’剛’ and’才’, ’做(什麼)’ and ’幹(什麼)’ . Time delay activated tokens are proposed tokens that are targets at a later step, where selecting them effectively skips some intermediate tokens. We observe that these tokens are much more likely to be present at the start of the sequence.
LLM probability estimation errors: LLM probability estimates are generally aligned with the logical coherence of a sequence. However, a major discrepancy arises with repeating sequences. Due to the in-context learning abilities of LLMs, they tend to significantly overestimate the likelihood of ever-repeating sequences. This could lead to mode collapse during the entire decoding process.
6 Conclusion
The Generative Fusion Decoding (GFD) framework represents a leap forward in fusing auto-regressive text models from heterogeneous token spaces. Our experiments confirm GFD’s effectiveness, particularly in ASR tasks. For tonal languages like Chinese, it achieves a mixed error rate of 8.83% on the NTUML2021 dataset, surpassing previous benchmarks. In context-aware recognition, GFD demonstrated its prowess by a 17.7% Word Error Rate Reduction (WERR) on the ATCO2 dataset when utilizing prompts to both ASR and LLM . For OCR tasks, GFD’s fusion with LLMs resulted in a notable 38.06% improvement in exact matches on the NAF-Long dataset, underscoring the framework’s ability to enhance recognition of long text sequences. These experimental outcomes highlight GFD’s robustness and adaptability, showcasing its generality to leverage the extensive training and contextual understanding of LLMs to improve text recognition across various domains and challenges. We hope that our framework will motivate further research into more fusion techniques that better leverages the existing capabilties of the pre-trained models.